PECULARITIES OF ISOTROPIC RAYLEIGH LIGHT SCATTERING IN AQUEOUS SOLUTIONS OF CAR-BOXYLIC ACIDS

\underline{S. A. OSMANOV}, Samarkand State University, 703004 Samarkand, Univ. Blvd, 15, Uzbekistan (email: barry@uni.uzsci.net).